# Evolution of a Single Impurity Across the Superfluid–Mott insulator Transition in the Bose–Hubbard Model

### v1

@@ V Impurity in a Mott-insulating bath
A key feature of the Mott-insulator state is that bath occupation can change only in integer units. This quantization strongly constrains the impurity response: under a repulsive impurity-bath coupling ($U_{\mathrm{ib}}/t>0$), the impurity may lower its energy only by expelling one bath boson from its vicinity; under an attractive impurity-bath coupling ($U_{\mathrm{ib}}/t<0$), it may recruit one additional boson. Extended deformations, multi-particle halos, and continuous dressing clouds are forbidden by the incompressibility of the bath. Thus, impurity localization in the MI proceeds via discrete $\pm 1$–particle defects rather than via compressibility-enabled screening.
In the following two subsections we analyze these two cases separately at fixed bath interaction $U_{\mathrm{b}}/t=20.0$, representative of the deep Mott-insulating regime. For repulsive impurity-bath coupling, we show that the impurity evolves from a weakly dressed, nearly-free defect at small $U_{\mathrm{ib}}/t$ to a fully localized vacancy defect in which exactly one bath boson is expelled. For attractive impurity-bath coupling, we find an analogous progression from a lightly dressed defect to a localized particle defect that binds one additional boson.
Together, these two trajectories establish the characteristic, quantized signatures of impurity localization in an incompressible bath and highlight their fundamental difference from the continuous, bath interaction-driven dressing found on the superfluid side.
Figure 6: Repulsive impurity in a Mott-insulating bath: cumulative bath-density deformation. (a) Cumulative bath-density deformation $\Delta N(R)$ for increasing repulsive coupling $U_{\mathrm{ib}}/t$ at $U_{\mathrm{b}}/t=20.0$. For weak $U_{\mathrm{ib}}/t$, the response shows only a shallow minimum, indicative of a nearly-free defect. For $U_{\mathrm{ib}}/t\!\gtrsim\!13.0$, the minimum approaches $\Delta N_{\min}\approx-1$, signaling the expulsion of a single bath particle. (b) Radius of the minimum, $R(\Delta N_{\min})$, which sharply increases at the localization threshold. (c) Magnitude of the minimum $|\Delta N_{\min}(R)|$, saturating to unity once the vacancy defect is formed.
Figure 7: Repulsive impurity in a Mott-insulating bath: radial impurity–centered correlator. Radical-averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ evolves from a shallow, short-range dip at weak $U_{\mathrm{ib}}/t$ (nearly-free defect) to a deep on-site depletion with a compensating positive overshoot at larger $R$ for $U_{\mathrm{ib}}/t\!\gtrsim\!12.0$, indicating formation of a saturated vacancy defect. Inset: on-site contrast $C_{\mathrm{ib}}(0)$, showing a sharp drop at the same threshold, consistent with the expulsion of one bath particle from the impurity site.
@@ V.1 Repulsive impurity in a Mott-insulating bath
We now investigate the behavior of a repulsive impurity ($U_{\mathrm{ib}}/t>0$) embedded in a deep Mott-insulating bath with $U_{\mathrm{b}}/t=20.0$. In this regime, the background is incompressible, long-wavelength density fluctuations are frozen, and only local particle–hole excitations are accessible. Under these conditions the impurity cannot form an extended polaronic cloud: instead, its dressing is strictly local, and localization occurs through discrete defect formation. Figures 68 summarize our analysis.
@@
Strong repulsion: formation of a localized vacancy defect.
@@
The vacancy-defect formation is accompanied by a collapse of coherent motion. The winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ drops rapidly once $U_{\mathrm{ib}}/t\simeq 13.0$ and approaches zero for larger impurity-bath couplings [Fig. 8(a)], signaling complete localization of the impurity–vacancy composite object. Simultaneously, the total bath particle number $N_{\mathrm{b}}$ decreases by exactly one particle [Fig. 8(b)], confirming that the impurity binds to a vacancy in the Mott-insulating background. This behavior is fundamentally distinct from the superfluid case discussed in companion: there, localization proceeds through a continuous redistribution of density, while in the Mott-insulating regime it is enforced by the discrete and gapped nature of particle–hole excitations. The impurity therefore transitions from a nearly-free defect to a fully localized vacancy defect as $U_{\mathrm{ib}}/t$ increases, providing a demonstration of defect-mediated localization in an incompressible quantum state.
Figure 8: Impurity coherence and bath particle number for repulsive impurity-bath coupling in the MI. (a) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$. Finite winding at small $U_{\mathrm{ib}}/t$ confirms coherent motion of a nearly-free defect, whereas a rapid collapse when $U_{\mathrm{ib}}/t>13.0$ marks localization of the impurity–vacancy composite. (b) Total bath occupation $N_{\mathrm{b}}$, showing a quantized decrease by exactly one particle once the vacancy defect is formed. This discrete change highlights that localization in the Mott-insulating regime proceeds by defect binding rather than extended density redistribution.
@@ V.2 Attractive impurity in a Mott-insulating bath
Figure 9: Attractive impurity in a Mott-insulating bath: cumulative bath-density deformation. (a) Cumulative bath-density excess $\Delta N(R)$ for increasing attraction $|U_{\mathrm{ib}}|/t$ at $U_{\mathbf{b}}/t=20.0$, showing the growth of a positive deformation near the impurity. (b) Radius of maximum, $R(\Delta N_{\max})$, extracted from panel (a). (c) Magnitude of maximum $\Delta N_{\max}(R)$. Its rapid increase to values $\gtrsim 2$ signals the recruitment of an integer number of bath bosons and the emergence of a localized particle-defect.
Figure 10: Attractive impurity in a Mott-insulating bath: radial impurity-centered correlator. Radially averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ for various $|U_{\mathrm{ib}}|/t$, showing the evolution from a weak local enhancement (nearly-free defect) to a strong, sharply localized peak (particle-defect). Inset: on-site contrast $C_{\mathrm{ib}}(0)$ versus $U_{\mathrm{ib}}/t$, illustrating the abrupt growth of local density once the particle-defect formation threshold is crossed.
@@
For $|U_{\mathrm{ib}}|/t\lesssim 13.0$, the impurity behaves as a nearly-free defect in the MI background. The cumulative bath response $\Delta N(R)$ [Fig. 9(a)] displays only a small positive bump, with a shallow maximum $\Delta N_{\max}\!\ll\!0.1$ located at $R(\Delta N_{\max})\approx 4$ [Fig. 9(b)]. This indicates weak, short-range density attraction toward the impurity. Consistently, the radial-averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ [Fig. 10] shows a modest on-site enhancement with negligible tail, and the on-site contrast $C_{\mathrm{ib}}(0)$ (inset) remains close to zero. The impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ remains finite and only slowly decreases with increasing $|U_{\mathrm{ib}}|/t$ [Fig. 11(a)], confirming that the impurity moves coherently through the rigid MI background. The total bath particle number $N_{\mathrm{b}}$ is pinned at its integer value [Fig. 11(b)], establishing that no bath particle has yet been trapped. Overall, the impurity forms only a weakly dressed defect, in sharp contrast to the extended droplets expected in a compressible superfluid environment.
Strong attraction: formation of a localized particle defect.
A qualitative change occurs around $|U_{\mathrm{ib}}|/t\simeq 13.0$$14.0$. The cumulative bath response develops a rapidly growing positive maximum $\Delta N_{\max}\sim 1$ [Fig. 9(c)], while its position $R(\Delta N_{\max})$ remains confined to the first a few shells, $R\approx 1$33 [Fig. 9(b)]. This indicates that the impurity abruptly binds an integer number of localized bath bosons—effectively one extra particle—into its vicinity, forming a compact particle-like defect. The clustering is evident in the local correlator: $C_{\mathrm{ib}}(0)$ increases sharply and reaches a large positive value [inset of Fig. 10], while the shape of $C_{\mathrm{ib}}(R)$ shows a strong on-site accumulation followed by a tiny compensating negative tail, which restores global particle-number conservation. Beyond this threshold, further increasing $|U_{\mathrm{ib}}|/t$ does not enlarge the accumulation cloud: the one-particle defect has fully saturated, and its spatial extent is limited by the on-site repulsion $U_{\mathrm{b}}/t$ of the Mott-insulating background.
Figure 11: Impurity coherence and bath particle number for attractive impurity-bath coupling in the MI. (a) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$, showing finite mobility at small $|U_{\mathrm{ib}}|/t$ (nearly-free defect) and a sharp suppression once a particle-defect forms. (b) Total bath particle number $N_{\mathrm{b}}$. The discrete jump by one particle near $|U_{\mathrm{ib}}|/t\!\approx\!13.0$–14.0 confirms the quantized recruitment of a bath boson and the formation of a localized composite defect.
@@
The formation of the particle-defect cluster is accompanied by a suppression of coherent impurity motion: $\langle W_{\mathrm{imp}}^{2}\rangle$ drops to zero for $|U_{\mathrm{ib}}|/t\gtrsim 14.0$ [Fig. 11(a)]. Meanwhile, the total bath occupation $N_{\mathrm{bath}}$ increases by exactly one particle [Fig. 11(b)], demonstrating that the impurity binds to an additional bath boson to form a localized composite defect. This mechanism sharply contrasts with the smooth droplet formation in a superfluid bath. In the MI, the lack of compressibility forbids extended halos; localization proceeds through a quantized recruitment of bath particles. The resulting localized object is thus best viewed as a particle defect, distinct from both polarons (SF) and vacancy defect (repulsive MI).
Together, these results establish two characteristic regimes for attractive impurities in an incompressible bath: (i) a weakly dressed nearly-free defect at small $|U_{\mathrm{ib}}|/t$; and (ii) a strongly bound, localized particle defect at larger $|U_{\mathrm{ib}}|/t$, marked by discrete particle capture and vanishing impurity winding.
Figure 12: Quasiparticle properties of an attractive impurity across the SF–MI transition at fixed impurity-bath coupling $U_{\mathrm{ib}}/t=-8.0$. Ground-state energy $E_{\mathrm{p}}/t$ (a), inverse effective mass $m^{*}/m_{0}$ (b), and quasiparticle weight $Z_{0}$ (c) as functions of bath interaction $U_{\mathrm{b}}/t$. The gray shaded band highlights the SF–MI crossover region ($U_{\mathrm{b}}/t\approx 16.5$$17.0$). In the superfluid regime, the impurity behaves as a mobile light polaron with finite dressing cloud and moderate mass enhancement. As the bath approaches the SF–MI transition, both $m^{*}/m_{0}$ and $Z_{0}$ approach their bare values, reflecting the compressibility-cotrolled collapse of dressing. In the deep Mott-insulating regime, the impurity becomes a nearly-free defect, without forming a localized bound defect or recruiting an extra bath boson. Unlike the repulsive case at $U_{\mathrm{ib}}/t=8.0$, no quantized defect or vacancy/polaron complex is formed.
@@ VI Crossing the superfluid–Mott insulator transition at fixed impurity–bath coupling
Figure 13: Real-space impurity dressing for attractive impurity across the SF–MI transition at fixed attractive impurity-bath coupling $U_{\mathrm{ib}}/t=-8.0$. (a) Cumulative bath-density deformation $\Delta N(R)$ for increasing $U_{\mathrm{b}}/t$. In the superfluid regime ($U_{\mathrm{b}}/t\lesssim 15.0$), the impurity generates an extended accumulation halo with maximum $\Delta N_{\mathrm{max}}\approx 0.5$ at $R\approx 4$, characteristic of a light polaron with finite spatial dressing. As the bath approaches the SF–MI critical region ($U_{\mathrm{b}}/t\sim 16.7$), both the amplitude and width of the halo collapse, reflecting the suppression of compressibility and the disappearance of long-range screening. Inside the Mott-insulating regime ($U_{\mathrm{b}}/t_{\mathrm{b}}\gtrsim 18.0$), $\Delta N(R)$ becomes almost flat with $\Delta N_{\mathrm{max}}\lesssim 0.02$ and $R(\Delta N_{\mathrm{max}})=1$, indicating a purely local, short-range distortion without real particle recruitment. (b) Radius of maximum $R(\Delta N_{\mathrm{max}})$ rapidly drops from $R\approx 4$ in the superfluid to $R=1$ in the Mott insulator phase, confirming that dressing becomes strictly local once compressibility vanishes. (c) Magnitude of maximum $\Delta N_{\mathrm{max}}$ with increasing $U_{\mathrm{b}}/t$. Its smooth decay and nearly zero value demonstrate that the impurity remains a weakly dressed, nearly–free defect, never reaching the particle defect regime for this impurity-bath coupling strength $U_{\mathrm{ib}}/t=-8.0$.
We now take a complementary viewpoint to the interaction-driven paths discussed above: instead of fixing the bath interaction $U_{\mathrm{b}}/t$ and varying impurity-bath coupling $U_{\mathrm{ib}}/t$, we perform “vertical cuts” in the $(U_{\mathrm{b}}/t,U_{\mathrm{ib}}/t)$ plane by scanning the bath interaction $U_{\mathrm{b}}/t$ while keeping the impurity–bath coupling $U_{\mathrm{ib}}/t$ fixed. This strategy directly reveals how impurity properties evolve as the bath medium crosses the superfluid–Mott insulator transition, and allows us to isolate the role of bath compressibility from that of the impurity-bath coupling strength itself.
We begin with the representative intermediate attractive impurity-bath coupling $U_{\mathrm{ib}}/t=-8.0$, which in a superfluid bath lies close to the self-trapping threshold identified in the phase diagram of our companion Letter [28]. This choice provides an ideal starting point to trace how the impurity’s coherence, mobility, and dressing cloud collapse as the bath loses compressibility: as $U_{\mathrm{b}}/t_{\mathrm{b}}$ is increased from the deep-superfluid regime ($U_{\mathrm{b}}/t=8.0$–16.0), through the critical region ($U_{\mathrm{b}}/t\simeq 16.7$), and into the Mott-insulating regime ($U_{\mathrm{b}}/t=17.0$–24.0), the impurity evolves from a well-defined light polaron in the SF to a nearly free defect in the MI.
@@
Along this compressibility-driven trajectory at fixed $U_{\mathrm{ib}}/t=-8.0$, we monitor both quasiparticle-level and bath-level quantities. On the impurity side we extract (i) the ground-state energy $E_{\mathrm{p}}$, (ii) the effective mass $m^{*}/m_{0}$ from the single-impurity Green’s function, (iii) the quasiparticle residue $Z_{0}$, and (iv) the impurity winding-number estimator $\langle W_{\mathrm{imp}}^{2}\rangle$, which directly quantifies the extent of coherent motion and serves as our primary marker of localization. On the bath side we measure (i) the superfluid density $\rho_{\mathrm{s}}$ from global winding statistics, which tracks macroscopic phase coherence and stiffness, and (ii) the radial-averaged impurity-centered density response encoded in $C_{\mathrm{ib}}(R)$ and its cumulative bath-density deformation $\Delta N(R)$, which reveal the range, sign, and degree of quantization of the dressing cloud. Taken together, these observables provide a unified microscopic picture linking global coherence, local density structure, and quasiparticle properties across the compressibility-controlled SF–MI transition.
@@
In contrast to the repulsive case analyzed in the companion Letter [28], where a moderate repulsion $U_{\mathrm{ib}}/t=+8.0$ drives the impurity into a fully localized defect in a deep Mott-insulating background. We now examine the attractive case at fixed impurity–bath coupling $U_{\mathrm{ib}}/t=-8.0$, and vary the bath interaction $U_{\mathrm{b}}/t$ across the superfluid–Mott insulator transition. This vertical trajectory is qualitatively different from the repulsive one: because the attraction is moderate and the bath becomes increasingly incompressible as $U_{\mathrm{b}}/t$ increases, the impurity can no longer sustain a heavy polaron nor form a bound cluster with multiple bath bosons. Instead, its dressing cloud continuously shrinks, and ultimately the impurity crosses over into a nearly-free defect even in the deep Mott-insulating regime. Importantly, at this coupling strength the impurity never becomes a localized defect and never recruits an additional bath particle; it remains mobile, with only minimal local density distortion.
Quasiparticle properties across the SF-MI transition.
@@
Figure 12 shows the evolution of the ground-state energy $E_{\mathrm{p}}/t$, inverse effective mass $m^{*}/m_{0}$, and quasiparticle residue $Z_{0}$ as functions of $U_{\mathrm{b}}/t$. In the superfluid regime ($U_{\mathrm{b}}/t\lesssim 16.0$), the impurity behaves as a light polaron characterized by extended density accumulation, moderately enhanced mass $m^{*}/m_{0}\simeq 1.1$$2.5$, and quasiparticle weight $Z_{0}\simeq 0.7$$0.9$. As the bath becomes less compressible and approaches the critical region ($U_{\mathrm{b}}/t\approx 16.7$, shaded band), both $m^{*}/m_{0}$ and $Z_{0}$ rapidly approach their bare values, reflecting a compressibility-controlled collapse of polaronic dressing. Deep in the Mott-insulating regime ($U_{\mathrm{b}}/t\geq 18.0$), bath density fluctuations are frozen, and the impurity behaves as a nearly-free defect with $m^{*}/m_{0}\approx 1$, $Z\to 1$, and only a small, strictly local density enhancement.
Bath responses across the SF-MI transition.

### v2

@@ V Impurity in a Mott-insulating bath
A key feature of the Mott-insulator state is that bath occupation can change only in integer units. This quantization strongly constrains the impurity response: under a repulsive impurity-bath coupling ($U_{\mathrm{ib}}/t>0$), the impurity may lower its energy only by expelling one bath boson from its vicinity; under an attractive impurity-bath coupling ($U_{\mathrm{ib}}/t<0$), it may recruit one additional boson. Extended deformations, multi-particle halos, and continuous dressing clouds are forbidden by the incompressibility of the bath. Thus, impurity self-trapping in the MI proceeds via discrete $\pm 1$–particle defects rather than via compressibility-enabled screening.
In the following two subsections we analyze these two cases separately at fixed bath interaction $U_{\mathrm{b}}/t=20.0$, representative of the deep Mott-insulating regime. For repulsive impurity-bath coupling, we show that the impurity evolves from a weakly dressed, nearly-free defect at small $U_{\mathrm{ib}}/t$ to a fully self-trapped vacancy defect in which exactly one bath boson is expelled. For attractive impurity-bath coupling, we find an analogous progression from a lightly dressed defect to a self-trapped particle defect that binds one additional boson.
Together, these two trajectories establish the characteristic, quantized signatures of impurity self-trapping in an incompressible bath and highlight their fundamental difference from the continuous, bath interaction-driven dressing found on the superfluid side.
Figure 6: Repulsive impurity in a Mott-insulating bath: cumulative bath-density deformation. (a) Cumulative bath-density deformation $\Delta N(R)$ for increasing repulsive coupling $U_{\mathrm{ib}}/t$ at $U_{\mathrm{b}}/t=20.0$. For weak $U_{\mathrm{ib}}/t$, the response shows only a shallow minimum, indicative of a nearly-free defect. For $U_{\mathrm{ib}}/t\!\gtrsim\!13.0$, the minimum approaches $\Delta N_{\min}\approx-1$, signaling the expulsion of a single bath particle. (b) Radius of the minimum, $R(\Delta N_{\min})$, which sharply increases at the self-trapping threshold. (c) Magnitude of the minimum $|\Delta N_{\min}(R)|$, saturating to unity once the vacancy defect is formed.
Figure 7: Repulsive impurity in a Mott-insulating bath: radial impurity–centered correlator. Radial-averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ evolves from a shallow, short-range dip at weak $U_{\mathrm{ib}}/t$ (nearly-free defect) to a deep on-site depletion for $U_{\mathrm{ib}}/t\!\gtrsim\!12.0$, indicating formation of a saturated vacancy defect. Inset: on-site contrast $C_{\mathrm{ib}}(0)$, showing a sharp drop at the same threshold, consistent with the expulsion of one bath particle from the impurity site.
@@ V.1 Repulsive impurity in a Mott-insulating bath
We now investigate the behavior of a repulsive impurity ($U_{\mathrm{ib}}/t>0$) embedded in a deep Mott-insulating bath with $U_{\mathrm{b}}/t=20.0$. In this regime, the background is incompressible, long-wavelength density fluctuations are frozen, and only local particle–hole excitations are accessible. Under these conditions the impurity cannot form an extended polaronic cloud: instead, its dressing is strictly local, and self-trapping occurs through discrete defect formation. Figures 68 summarize our analysis.
@@
Strong repulsion: formation of a self-trapped vacancy defect.
@@
The vacancy-defect formation is accompanied by a collapse of coherent motion. The winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ drops rapidly once $U_{\mathrm{ib}}/t\simeq 13.0$ and approaches zero for larger impurity-bath couplings [Fig. 8(a)], signaling complete self-trapping of the impurity–vacancy composite object. Simultaneously, the total bath particle number $N_{\mathrm{b}}$ decreases by exactly one particle [Fig. 8(b)], confirming that the impurity binds to a vacancy in the Mott-insulating background. This behavior is fundamentally distinct from the superfluid case discussed in companion: there, self-trapping proceeds through a continuous redistribution of density, while in the Mott-insulating regime it is enforced by the discrete and gapped nature of particle–hole excitations. The impurity therefore transitions from a nearly-free defect to a fully self-trapped vacancy defect as $U_{\mathrm{ib}}/t$ increases, providing a demonstration of defect-mediated self-trapping in an incompressible quantum state.
Figure 8: Impurity coherence and bath particle number for repulsive impurity-bath coupling in the MI. (a) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$. Finite winding at small $U_{\mathrm{ib}}/t$ confirms coherent motion of a nearly-free defect, whereas a rapid collapse when $U_{\mathrm{ib}}/t>13.0$ marks self-trapping of the impurity–vacancy composite. (b) Total bath occupation $N_{\mathrm{b}}$, showing a quantized decrease by exactly one particle once the vacancy defect is formed. This discrete change highlights that self-trapping in the Mott-insulating regime proceeds by defect binding rather than extended density redistribution.
@@ V.2 Attractive impurity in a Mott-insulating bath
Figure 9: Attractive impurity in a Mott-insulating bath: cumulative bath-density deformation. (a) Cumulative bath-density excess $\Delta N(R)$ for increasing attraction $|U_{\mathrm{ib}}|/t$ at $U_{\rm{b}}/t=20.0$, showing the growth of a positive deformation near the impurity. (b) Radius of maximum, $R(\Delta N_{\max})$, extracted from panel (a). (c) Magnitude of maximum $\Delta N_{\max}(R)$, saturating to unity once the particle defect is formed.
Figure 10: Attractive impurity in a Mott-insulating bath: radial impurity-centered correlator. Radially averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ for various $|U_{\mathrm{ib}}|/t$, showing the evolution from a weak local enhancement (nearly-free defect) to a strong, sharply self-trapped peak (particle-defect). Inset: on-site contrast $C_{\mathrm{ib}}(0)$ versus $U_{\mathrm{ib}}/t$, illustrating the abrupt growth of local density once the particle-defect formation threshold is crossed.
@@
For $|U_{\mathrm{ib}}|/t\lesssim 13.0$, the impurity behaves as a nearly-free defect in the MI background. The cumulative bath response $\Delta N(R)$ [Fig. 9(a)] displays only a small positive bump, with a shallow maximum $\Delta N_{\max}\!\ll\!0.1$ located at $R(\Delta N_{\max})\approx 4$ [Fig. 9(b)]. This indicates weak, short-range density attraction toward the impurity. Consistently, the radial-averaged impurity-centered correlator $C_{\mathrm{ib}}(R)$ [Fig. 10] shows a modest on-site enhancement with negligible tail, and the on-site contrast $C_{\mathrm{ib}}(0)$ (inset) remains close to zero. The impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$ remains finite and only slowly decreases with increasing $|U_{\mathrm{ib}}|/t$ [Fig. 11(a)], confirming that the impurity moves coherently through the rigid MI background. The total bath particle number $N_{\mathrm{b}}$ is pinned at its integer value [Fig. 11(b)], establishing that no bath particle has yet been self-trapped. Overall, the impurity forms only a weakly dressed defect, in sharp contrast to the extended droplets expected in a compressible superfluid environment.
Strong attraction: formation of a self-trapped particle defect.
A qualitative change occurs around $|U_{\mathrm{ib}}|/t\simeq 13.0$$14.0$. The cumulative bath response develops a rapidly growing positive maximum $\Delta N_{\max}\sim 1$ [Fig. 9(c)], while its position $R(\Delta N_{\max})$ remains confined to the first a few shells, $R\approx 1$33 [Fig. 9(b)]. This indicates that the impurity abruptly captures one additional bath boson, i.e., a quantized +1+1 change in the bath occupation, forming a compact particle-like defect. The clustering is evident in the local correlator: $C_{\mathrm{ib}}(0)$ increases sharply and reaches a large positive value [inset of Fig. 10], while the shape of $C_{\mathrm{ib}}(R)$ shows a strong on-site accumulation followed by a tiny tail. Beyond this threshold $|U_{\rm ib}|/t>14,0$, further increasing $|U_{\mathrm{ib}}|/t$ does not enlarge the accumulation cloud, only deepens the onsite $C_{\rm in}(0)$.
Figure 11: Impurity coherence and bath particle number for attractive impurity-bath coupling in the MI. (a) Impurity winding number $\langle W_{\mathrm{imp}}^{2}\rangle$, showing finite mobility at small $|U_{\mathrm{ib}}|/t$ (nearly-free defect) and a sharp suppression once a particle-defect forms. (b) Total bath particle number $N_{\mathrm{b}}$. The discrete jump by one particle near $|U_{\mathrm{ib}}|/t\!\approx\!13.0$–14.0 confirms the quantized recruitment of a bath boson and the formation of a self-trapped composite defect.
@@
The formation of the particle-defect cluster is accompanied by a suppression of coherent impurity motion: $\langle W_{\mathrm{imp}}^{2}\rangle$ drops to zero for $|U_{\mathrm{ib}}|/t\gtrsim 14.0$ [Fig. 11(a)]. Meanwhile, the total bath occupation $N_{\mathrm{b}}$ increases by exactly one particle [Fig. 11(b)], demonstrating that the impurity binds to an additional bath boson to form a self-trapped composite defect. This mechanism sharply contrasts with the smooth droplet formation in a superfluid bath. In the MI, the lack of compressibility forbids extended halos; self-trapping proceeds through a quantized recruitment of bath particles. The resulting self-trapped object is thus best viewed as a particle defect, distinct from both polarons (SF) and vacancy defect (repulsive MI).
Together, these results establish two characteristic regimes for attractive impurities in an incompressible bath: (i) a weakly dressed nearly-free defect at small $|U_{\mathrm{ib}}|/t$; and (ii) a strongly bound, self-trapped particle defect at larger $|U_{\mathrm{ib}}|/t$, marked by discrete particle capture and vanishing impurity winding.
Figure 12: Quasiparticle properties of an attractive impurity across the SF–MI transition at fixed impurity-bath coupling $U_{\mathrm{ib}}/t=-8.0$. Ground-state energy $E_{\mathrm{p}}/t$ (a), effective mass $m^{*}/m_{0}$ (b), and quasiparticle weight $Z_{0}$ (c) as functions of bath interaction $U_{\mathrm{b}}/t$. The gray shaded band highlights the SF–MI crossover region ($U_{\mathrm{b}}/t\approx 16.5$$17.0$). In the superfluid regime, the impurity behaves as a mobile light polaron with finite dressing cloud and moderate mass enhancement. As the bath approaches the SF–MI transition, both $m^{*}/m_{0}$ and $Z_{0}$ approach their bare values, reflecting the compressibility-cotrolled collapse of dressing. In the deep Mott-insulating regime, the impurity becomes a nearly-free defect, without forming a self-trapped bound defect or recruiting an extra bath boson. Unlike the repulsive case at $U_{\mathrm{ib}}/t=8.0$, no quantized defect or vacancy/polaron complex is formed.
@@ VI Crossing the superfluid–Mott insulator transition at fixed impurity–bath coupling
Figure 13: Real-space impurity dressing for an attractive impurity across the SF–MI transition at fixed impurity–bath coupling $U_{\mathrm{ib}}/t=-8.0$. (a) Cumulative bath-density deformation $\Delta N(R)$ for increasing $U_{\mathrm{b}}/t$. In the superfluid regime ($U_{\mathrm{b}}/t\lesssim 15.0$), the impurity generates an extended accumulation halo with maximum $\Delta N_{\mathrm{max}}\approx 0.5$ at $R\approx 4$, characteristic of a light polaron with finite spatial dressing. As the bath approaches the SF–MI critical region ($U_{\mathrm{b}}/t\sim 16.7$), both the amplitude and width of the halo collapse, reflecting the suppression of bath compressibility and the disappearance of long-range screening. Deep in the Mott-insulating regime ($U_{\mathrm{b}}/t\gtrsim 18.0$), $\Delta N(R)$ becomes almost flat with $\Delta N_{\mathrm{max}}\lesssim 0.02$ and $R(\Delta N_{\mathrm{max}})=1$, indicating a purely local, short-range distortion without quantized particle recruitment. (b) Radius of maximum $R(\Delta N_{\mathrm{max}})$ rapidly drops from $R\approx 4$ in the superfluid to $R=1$ in the Mott-insulating phase, confirming that dressing becomes strictly local once compressibility vanishes. (c) Magnitude of maximum $\Delta N_{\mathrm{max}}$ with increasing $U_{\mathrm{b}}/t$. Its smooth decay and nearly zero value demonstrate that the impurity remains a weakly dressed, nearly free defect, never reaching the particle-defect regime for this impurity–bath coupling strength $U_{\mathrm{ib}}/t=-8.0$.
We now take a complementary viewpoint to the interaction-driven paths discussed above: instead of fixing the bath interaction $U_{\mathrm{b}}/t$ and varying impurity–bath coupling $U_{\mathrm{ib}}/t$, we perform “vertical cuts” in the $(U_{\mathrm{b}}/t,U_{\mathrm{ib}}/t)$ plane by scanning the bath interaction $U_{\mathrm{b}}/t$ while keeping the impurity–bath coupling $U_{\mathrm{ib}}/t$ fixed. This strategy directly reveals how impurity properties evolve as the bath medium crosses the superfluid–Mott-insulator transition, and allows us to isolate the role of bath compressibility from that of the impurity–bath coupling strength itself.
We begin with the representative intermediate attractive impurity–bath coupling $U_{\mathrm{ib}}/t=-8.0$, which in a superfluid bath lies close to the self-trapping threshold between the light polaron region and heavy polaron region identified in the phase diagram of our companion Letter [28]. This choice provides an ideal starting point to trace how the impurity’s coherence, mobility, and dressing cloud collapse as the bath loses compressibility: as $U_{\mathrm{b}}/t$ is increased from the deep-superfluid regime ($U_{\mathrm{b}}/t=8.0$–16.0), through the critical region ($U_{\mathrm{b}}/t\simeq 16.7$), and into the Mott-insulating regime ($U_{\mathrm{b}}/t=17.0$–24.0), the impurity evolves from a well-defined light polaron in the SF to a nearly free defect in the MI.
@@
Along this compressibility-driven trajectory at fixed $U_{\mathrm{ib}}/t=-8.0$, we monitor both quasiparticle-level and bath-level quantities. On the impurity side we extract (i) the ground-state energy $E_{\mathrm{p}}$, (ii) the effective mass $m^{*}/m_{0}$ from the single-impurity Green’s function, (iii) the quasiparticle residue $Z_{0}$, and (iv) the impurity winding-number estimator $\langle W_{\mathrm{imp}}^{2}\rangle$. On the bath side we measure (i) the superfluid density $\rho_{\mathrm{b}}$ from global winding statistics, which tracks macroscopic phase coherence and stiffness, and (ii) the radial-averaged impurity-centered density response encoded in $C_{\mathrm{ib}}(R)$ and its cumulative bath-density deformation $\Delta N(R)$, which reveal the range and amplitude of the dressing cloud. Taken together, these observables provide a unified microscopic picture linking global coherence, local density structure, and quasiparticle properties across the compressibility-controlled SF–MI transition.
@@
In contrast to the repulsive case analyzed in the companion Letter [28], where a moderate repulsion $U_{\mathrm{ib}}/t=+8.0$ drives the impurity into a self-trapped defect in a deep Mott-insulating background, we now examine the attractive case at fixed impurity–bath coupling $U_{\mathrm{ib}}/t=-8.0$, and vary the bath interaction $U_{\mathrm{b}}/t$ across the superfluid–Mott-insulator transition. This vertical trajectory is qualitatively different from the repulsive one: because the attraction is moderate and the bath becomes increasingly incompressible as $U_{\mathrm{b}}/t$ increases, the impurity can no longer sustain a heavy polaron nor form a bound cluster with multiple bath bosons. Instead, its dressing cloud continuously shrinks, and ultimately the impurity crosses over into a nearly free defect even in the deep Mott-insulating regime. Importantly, at this coupling strength the impurity never binds into a particle defect and never recruits an additional bath particle; it remains mobile, with only minimal local density distortion.
Quasiparticle properties across the SF–MI transition.
@@
Figure 12 shows the evolution of the ground-state energy $E_{\mathrm{p}}/t$, effective mass $m^{*}/m_{0}$, and quasiparticle residue $Z_{0}$ as functions of $U_{\mathrm{b}}/t$. In the superfluid regime ($U_{\mathrm{b}}/t\lesssim 16.0$), the impurity behaves as a light polaron characterized by extended density accumulation, moderately enhanced mass $m^{*}/m_{0}\simeq 1.1$$2.5$, and quasiparticle weight $Z_{0}\simeq 0.7$$0.9$. As the bath becomes less compressible and approaches the critical region ($U_{\mathrm{b}}/t\approx 16.7$, shaded band), both $m^{*}/m_{0}$ and $Z_{0}$ rapidly approach their bare values, reflecting a compressibility-controlled collapse of polaronic dressing. Deep in the Mott-insulating regime ($U_{\mathrm{b}}/t\geq 18.0$), bath density fluctuations are frozen, and the impurity behaves as a nearly free defect with $m^{*}/m_{0}\approx 1$, $Z_{0}\to 1$, and only a small, strictly local density enhancement.
Bath responses across the SF–MI transition.